# Advances in the Shannon Capacity of Graphs

### v2

@@ 2.4 Concluding preliminaries
We next provide a useful lower bound for the fractional independence number and the fractional chromatic number (see [37, Proposition 3.1.1]).
@@
Let $\mathsf{G}$ be a simple graph of order $n$, then
@@
In continuation to Definition 2.19, we next provide some known invariants of Kneser graphs.
@@
For $\mathsf{G}$ to be a universal graph, the equality $\alpha(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})$ has to hold (see Theorem 2.6), thus, $\mathsf{G}$ is universal if and only if
@@
The Petersen graph $\mathsf{G}=\operatorname{\mathsf{K}}(5,2)$ satisfies $\alpha(\mathsf{G})=\Theta(\mathsf{G})$, yet it is not universal by Corollary 2.1.
@@ 3 The Shannon capacity of polynomials in graphs
We note that the simple counterexample constructed earlier by Haemers [15] plays a key role in this argument and was subsequently employed by Alon [7] in establishing (66).
@@
Let $\mathbb{N}[x_{1},\ldots,x_{\ell}]$ be the set of nonzero polynomials in the variables $x_{1},\ldots,x_{\ell}$ with nonnegative integer coefficients. A polynomial in graphs, $p(\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell})$, is defined by interpreting graph addition as disjoint union and graph multiplication as strong product.
@@
Let $\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell}$ be simple graphs, with $\ell\in\mathbb{N}$. Then for every $p\in\mathbb{N}[x_{1},\ldots,x_{\ell}]$,
@@
If $\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell}$ are all Kneser graphs or self-complementary vertex-transitive graphs, then by Theorem 3.2, equality (68) holds for every polynomial $p\in\mathbb{N}[x_{1},\ldots,x_{\ell}]$. In particular, if $\mathsf{G}_{i}=\operatorname{\mathsf{K}}(n_{i},r_{i})$ (with $n_{i}\geq 2r_{i}$) for every $i\in[\ell]$, then by Theorem 2.23
@@
Similarly, if $\mathsf{G}_{i}$ is self-complementary and vertex-transitive for all $i\in[\ell]$, then by [39, Theorem 3.26]
@@
Next we give second sufficient conditions for equality (68) to hold.
@@
It is worth noting that our sufficient conditions for equality (68) to hold are less restrictive than those hinted in Shannon’s paper [1], which are next presented. The following theorem is derived directly from Lemma 3.2 and the sandwich theorem. We provide the original proof in Appendix C.
@@
Indeed, the invariants of the Kneser graph and its complement are known, and by Theorem 2.23 (since $\alpha(\mathsf{G})=\operatorname{\omega}(\overline{\mathsf{G}})$, where $\overline{\mathsf{G}}=\operatorname{\mathsf{K}}(n,r)$ with $r$ and $n$ as above)
@@
Since the independence number and the fractional independence number coincide and are both equal to $\frac{n}{r}$ then $\Theta(\mathsf{G})=\frac{n}{r}$. Hence, the assumption $n>2r$ implies that
@@
so the condition $\Theta(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})$ implies that the equality $\Theta(\mathsf{G})=\vartheta(\mathsf{G})$ holds. Hence, Theorem 3.3 imposes a stronger condition than the one in Theorem 3.2 on $\ell-1$ of the graphs, while no condition is imposed on the $\ell$-th graph.
@@
$\alpha_{\mathrm{f}}(\mathsf{G}_{i})=\frac{\binom{n_{i}}{r_{i}}}{\lfloor\frac{n_{i}}{r_{i}}\rfloor},\quad\forall\,i\in[\ell].$ (93)
@@ 4 The Shannon capacity of Tadpole graphs
Next, if $k>3$, then the clique number of $\mathrm{T}({k},{\ell})$ is $2$ (since there are no triangles), thus, every clique in $\mathrm{T}({k},{\ell})$ is either an edge or a single vertex. In this case, if $k$ is even, it is possible to cover the cycle $\operatorname{\mathsf{C}}_{k}$ by $\frac{k}{2}$ edges (cliques) and the path can be covered by $\left\lceil\frac{\ell}{2}\right\rceil$ edges, and thus $\operatorname{\sigma}(\mathrm{T}({k},{\ell}))=\frac{k}{2}+\left\lceil\frac{\ell}{2}\right\rceil$. And, if $k$ is odd and $\ell$ is odd, then one can cover $\mathrm{T}({k},{\ell})$ by $\frac{k+\ell}{2}$ edges (by covering $\operatorname{\mathsf{C}}_{k}$ by $\frac{k-1}{2}$ edges, excluding the vertex of $\operatorname{\mathsf{C}}_{k}$ that is adjacent to a pendant vertex of $\operatorname{\mathsf{P}}_{\ell}$, and then covering the remaining vertices of $\mathrm{T}({k},{\ell})$ by $\frac{\ell+1}{2}$ edges). Overall, in both instances we got
@@
which gives (by using again the sandwich theorem in (50) and Lemma 4.1),
@@
Finally, assume that $k\geq 5$ is odd and $\ell\geq 2$ is even (if $\ell=0$ then $\mathrm{T}({k},{0})=\operatorname{\mathsf{C}}_{k}$ and the result follows from Theorem 2.25). By deleting the edge that connects the cycle $\operatorname{\mathsf{C}}_{k}$ with the path $\operatorname{\mathsf{P}}_{\ell}$, it follows that $\operatorname{\mathsf{C}}_{k}+\operatorname{\mathsf{P}}_{\ell}$ is a spanning subgraph of $\mathrm{T}({k},{\ell})$. On the other hand, by deleting the leftmost vertex of $\operatorname{\mathsf{P}}_{\ell}$ (i.e., the vertex that is of distance $1$ from the cycle $\operatorname{\mathsf{C}}_{k}$), and the two edges that are incident to it in $\mathrm{T}({k},{\ell})$, it follows that $\operatorname{\mathsf{C}}_{k}+\operatorname{\mathsf{P}}_{\ell-1}$ is an induced subgraph of $\mathrm{T}({k},{\ell})$. Hence, by Theorem 2.18,
@@
First, if $k=3$ or $k\geq 4$ is even, or $k\geq 5$ is odd and $\ell\geq 1$ is odd, then by Lemmas 4.1 and 4.2,
@@
Let $\mathsf{G}_{1}=\mathrm{T}({5},{6})$, then by Corollary 4.1,
@@
The upper bound was proved in Corollary 4.1. We next prove the improved lower bound. By Hales’ result [32, Theorem 7.1], if $k\geq 3$ is odd, then
@@
$\left(\frac{k}{1+\sec\frac{\pi}{k}}+\frac{\ell}{2}\right)-\left(\sqrt{\left(\frac{k-1}{2}\right)^{2}+\left\lfloor\frac{k-1}{4}\right\rfloor}+\frac{\ell}{2}\right)$
@@
Theorem 4.2 gives an improved lower bound that relies on a lower bound on the Shannon capacity of odd cycles that was constructed by Hales (see [32]). Since Hales’ result, better lower bounds for odd cycles were constructed, and can be used to improve the lower bound in special cases. For better lower bounds on higher powers of odd cycles see [27, 41]. The next theorem shows an improvement to the bound in (124), for special cases using a result from [27].
@@
The proof of Theorem 4.3 relies on the lower bound on $\alpha(\operatorname{\mathsf{C}}_{k}^{d})$ given in (4), as presented in [27, Theorem 1.4]. It is worth noting that in [27], Bohman conjectured that the lower bound is the exact value of the independence number (see [27, Conjecture 1.5]). In a subsequent paper by Bohman, Holzman and Natarajan [41], the conjecture was confirmed for a countably infinite subset of these values. In particular, by substituting $d=3$, which gives $k=8n+5$, it was proved in [41] that if $8n+5$ is a prime number, then the lower bound coincides with the exact value of the independence number. Specifically, by [41, Theorem 1], the independence number of the strong cubed power of the cycle graph $\operatorname{\mathsf{C}}_{8n+5}$ is given by
@@ 5 When the graph capacity is not attained by the independence number of any finite power?
This problem was explored by Guo and Watanabe [10], and a family of disconnected graphs that satisfies (134) was constructed. In this section, we start by presenting the method from [10], and then we provide two original approaches. The first original approach relies on a Dedekind’s lemma in number theory (1858), and it uses a similar concept of proof to the one in Example 5.1. The second original approach uses the result from [10] to construct a countably infinite family of connected graphs whose capacity is strictly larger than the independence number of any finite (strong) power of the graph.
@@
It therefore follows that, for all $k\in\mathbb{N}$, $\Theta(\mathsf{G})^{k}\neq\alpha(\mathsf{G}^{k})$ since the independence number is an integer. Consequently, (134) holds.
@@
We next consider three approaches for the construction of families of graphs whose Shannon capacities satisfies the condition in (134). The first approach is due to Guo and Watanabe [10] with a simplified proof (the original proof from [10] is provided in Appendix D), and the second and third approaches of constructing such graph families are original.
@@ 5.1 First approach
where the last equality holds by Lemma 2.2. Next, by the assumption that $\Theta(\mathsf{H})>\alpha(\mathsf{H})$ and since $\Theta(\mathsf{H}^{m})=\Theta(\mathsf{H})^{m}\geq\alpha(\mathsf{H})^{m}$ for all $m\in\mathbb{N}$ with strict inequality if $m=1$ (by assumption), it follows that for all $k\in\mathbb{N}$
@@
$=(\Theta(\mathsf{G})+\Theta(\mathsf{H}))^{k}.$ (140)
@@
$\Theta(\operatorname{\mathsf{C}}_{2k+1})$ $\geq\sqrt{\alpha(\operatorname{\mathsf{C}}_{2k+1}^{2})}$
$=\sqrt{k^{2}+\bigg\lfloor\frac{k}{2}\bigg\rfloor}$
$>k$
$=\alpha(\operatorname{\mathsf{C}}_{2k+1}).$ (144)
@@ 5.2 Second approach
Let $m\in\mathbb{N}$ be a natural number such that $\sqrt{m}\in\mathbb{Q}$. Let $n_{0}\in\mathbb{N}$ be the smallest natural number such that $n_{0}\sqrt{m}\in\mathbb{N}$. If $\sqrt{m}\notin\mathbb{N}$, then there exists $\ell\in\mathbb{N}$ such that $0<\sqrt{m}-\ell<1$. Setting $n_{1}=n_{0}(\sqrt{m}-\ell)$, we get $n_{1}\in\mathbb{N}$ and $0<n_{1}<n_{0}$. Also $n_{1}\sqrt{m}=n_{0}m-\ell(n_{0}\sqrt{m})\in\mathbb{N}$. This leads to a contradiction to the choice of $n_{0}$. ∎
@@
Let $r\geq 2$, and let $\mathsf{G}_{1},\mathsf{G}_{2},\ldots,\mathsf{G}_{r}$ be graphs such that, for every $\ell\in[r]$,
@@
    The graph $\mathsf{G}_{\ell}$ is either a Kneser graph or self-complementary strongly regular or self-complementary vertex-transitive.
@@
    There exists a single $\ell_{0}\in[r]$ such that $\mathsf{G}_{\ell_{0}}$ is self-complementary vertex-transitive or self-complementary strongly regular on $n_{\ell_{0}}$ vertices, where $n_{\ell_{0}}$ is not a square of an integer.
Let $p\in\mathbb{N}[x_{1},\ldots,x_{r}]$ be a polynomial whose nonzero coefficients are natural numbers such that
@@
    The variable $x_{\ell_{0}}$ doesn’t occur in (at least) one of the monomials in $p(x_{1},\ldots,x_{r})$.
@@
For all $\ell\in[r]$, $\Theta(\mathsf{G}_{\ell})=\vartheta(\mathsf{G}_{\ell})$ (this follows from the first assumption on $\mathsf{G}_{\ell}$). Thus, by Corollary 3.1
@@
Let $r\geq 2$, and let $\mathsf{G}_{1},\mathsf{G}_{2},\ldots,\mathsf{G}_{r}$ be graphs such that, for every $\ell\in[r]$,
@@
Next we show another example, this time using Paley graphs.
@@
is not attained by the independence number of any finite strong power of $\mathsf{G}$. This follows from Corollary 5.2, as the Paley graphs are self-complementary and strongly regular, and the restrictions on $q_{1},\ldots,q_{\ell}$ guaranty that the conditions of Corollary 5.2 hold.

### v3

@@ 2.4 Concluding preliminaries
We provide a useful lower bound for the fractional independence number and the fractional chromatic number (see [37, Proposition 3.1.1]).
@@
Let $\mathsf{G}$ be a simple graph of order $n$. Then,
@@
The following sandwich theorem for the Lovász $\vartheta$-function is remarkable in that it provides a computable bound on graph invariants that are NP-hard to compute.
@@
In continuation to Definition 2.19, we provide some known invariants of Kneser graphs.
@@
For $\mathsf{G}$ to be a universal graph, the equality $\alpha(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})$ should hold (see Theorem 2.6). Thus, $\mathsf{G}$ is universal if and only if
@@
The Petersen graph $\mathsf{G}=\operatorname{\mathsf{K}}(5,2)$ satisfies $\alpha(\mathsf{G})=4=\Theta(\mathsf{G})$, yet it is not universal by Corollary 2.1.
@@ 3 The Shannon capacity of polynomials in graphs
We note that the simple counterexample constructed by Haemers [15] plays a key role in this argument and was later used by Alon [7] to establish (66).
@@
Let $\mathbb{N}[x_{1},\ldots,x_{\ell}]$ be the set of nonzero polynomials in the variables $x_{1},\ldots,x_{\ell}$ with nonnegative integer coefficients. A polynomial in graphs, $p(\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell})$, is defined by interpreting graph addition as the disjoint union and graph multiplication as the strong product.
@@
Let $\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell}$ be simple graphs, with $\ell\in\mathbb{N}$. Then, for every $p\in\mathbb{N}[x_{1},\ldots,x_{\ell}]$,
@@
If $\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell}$ are all Kneser graphs or self-complementary vertex-transitive graphs, then by Theorem 3.2, equality (68) holds for every polynomial $p\in\mathbb{N}[x_{1},\ldots,x_{\ell}]$. In particular, if $\mathsf{G}_{i}=\operatorname{\mathsf{K}}(n_{i},r_{i})$ (with $n_{i}\geq 2r_{i}$) for every $i\in[\ell]$, then by Theorem 2.23,
@@
Similarly, if $\mathsf{G}_{i}$ is a self-complementary and vertex-transitive graph for all $i\in[\ell]$, then by [39, Theorem 3.26],
@@
Next, we give a second sufficient condition for equality (68) to hold.
@@
It is worth noting that our sufficient conditions for equality (68) to hold are less restrictive than those hinted at in Shannon’s paper [1], which are presented next. The following theorem is derived directly from Lemma 3.2 and the sandwich theorem. We provide the original proof in Appendix C.
@@
Indeed, the invariants of the Kneser graph and its complement are known, and by Theorem 2.23 (since $\alpha(\mathsf{G})=\operatorname{\omega}(\overline{\mathsf{G}})$, where $\overline{\mathsf{G}}=\operatorname{\mathsf{K}}(n,r)$ with $r$ and $n$ as above),
@@
Since the independence number and the fractional independence number coincide and are both equal to $\frac{n}{r}$, then $\Theta(\mathsf{G})=\frac{n}{r}$. Hence, the assumption $n>2r$ implies that
@@
[…]
@@
$\alpha_{\mathrm{f}}(\mathsf{G}_{i})=\frac{\binom{n_{i}}{r_{i}}}{\bigl\lfloor\frac{n_{i}}{r_{i}}\bigr\rfloor},\quad\forall\,i\in[\ell].$ (93)
@@ 4 The Shannon capacity of Tadpole graphs
Next, if $k>3$, then the clique number of $\mathrm{T}({k},{\ell})$ is $2$ (since there are no triangles), thus every clique in $\mathrm{T}({k},{\ell})$ is either an edge or a single vertex. In this case, if $k$ is even, it is possible to cover the cycle $\operatorname{\mathsf{C}}_{k}$ by $\frac{k}{2}$ edges (cliques), the path $\operatorname{\mathsf{P}}_{\ell}$ can be covered by $\left\lceil\frac{\ell}{2}\right\rceil$ edges, and thus $\operatorname{\sigma}(\mathrm{T}({k},{\ell}))=\frac{k}{2}+\left\lceil\frac{\ell}{2}\right\rceil$. Moreover, if $k$ and $\ell$ are odd, then one can cover $\mathrm{T}({k},{\ell})$ by $\frac{k+\ell}{2}$ edges (by covering $\operatorname{\mathsf{C}}_{k}$ by $\frac{k-1}{2}$ edges, excluding the vertex of $\operatorname{\mathsf{C}}_{k}$ that is adjacent to a pendant vertex of $\operatorname{\mathsf{P}}_{\ell}$, and then covering the remaining vertices of $\mathrm{T}({k},{\ell})$ by $\frac{\ell+1}{2}$ edges). Overall, in both instances we obtain
@@
which gives, by again using the sandwich theorem in (50) and Lemma 4.1,
@@
Finally, assume that $k\geq 5$ is odd and $\ell\geq 2$ is even (if $\ell=0$, then $\mathrm{T}({k},{0})=\operatorname{\mathsf{C}}_{k}$, and the result follows from Theorem 2.25). By deleting the edge that connects the cycle $\operatorname{\mathsf{C}}_{k}$ with the path $\operatorname{\mathsf{P}}_{\ell}$, it follows that $\operatorname{\mathsf{C}}_{k}+\operatorname{\mathsf{P}}_{\ell}$ is a spanning subgraph of $\mathrm{T}({k},{\ell})$. On the other hand, by deleting the leftmost vertex of $\operatorname{\mathsf{P}}_{\ell}$ (i.e., the vertex that is of distance $1$ from the cycle $\operatorname{\mathsf{C}}_{k}$) and the two edges that are incident to it in $\mathrm{T}({k},{\ell})$, it follows that $\operatorname{\mathsf{C}}_{k}+\operatorname{\mathsf{P}}_{\ell-1}$ is an induced subgraph of $\mathrm{T}({k},{\ell})$. Hence, by Theorem 2.18,
@@
First, if $k=3$, $k\geq 4$ is even, or $k\geq 5$ is odd and $\ell\geq 1$ is odd, then by Lemmas 4.1 and 4.2,
@@
Let $\mathsf{G}_{1}=\mathrm{T}({5},{6})$. Then, by Corollary 4.1,
@@
The upper bound was proved in Corollary 4.1. Next, we prove the improved lower bound. By Hales’ result [32, Theorem 7.1], if $k\geq 3$ is odd, then
@@
$\hskip-5.69046pt\left(\frac{k}{1+\sec\frac{\pi}{k}}+\frac{\ell}{2}\right)-\left(\sqrt{\left(\frac{k-1}{2}\right)^{2}+\left\lfloor\frac{k-1}{4}\right\rfloor}+\frac{\ell}{2}\right)$
@@
Theorem 4.2 gives an improved lower bound that relies on a lower bound on the Shannon capacity of odd cycles that was constructed by Hales (see [32]). Since Hales’ result, better lower bounds for odd cycles were constructed, and can be used to improve the lower bound in special cases. For better lower bounds on higher powers of odd cycles see [27, 41]. The next theorem shows an improvement to the bound in (124) for special cases using a result from [27].
@@
The proof of Theorem 4.3 relies on the lower bound on $\alpha(\operatorname{\mathsf{C}}_{k}^{d})$ given in (4), as presented in [27, Theorem 1.4]. It is worth noting that in [27], Bohman conjectured that the lower bound is the exact value of the independence number (see [27, Conjecture 1.5]). In a subsequent paper by Bohman, Holzman, and Natarajan [41], the conjecture was confirmed for a countably infinite subset of these values. In particular, by substituting $d=3$, which gives $k=8n+5$, it was proved in [41] that if $8n+5$ is a prime number, then the lower bound coincides with the exact value of the independence number. Specifically, by [41, Theorem 1], the independence number of the strong cubed power of the cycle graph $\operatorname{\mathsf{C}}_{8n+5}$ is given by
@@ 5 When the graph capacity is not attained by the independence number of any finite power?
This problem was explored by Guo and Watanabe [10], and a family of disconnected graphs that satisfies (134) was constructed. In this section, we start by presenting the method from [10], and then we provide two original approaches. The first original approach relies on Dedekind’s lemma in number theory (see Lemma 5.1), and it uses a similar concept of proof to the one in Example 5.1. The second original approach uses the result from [10] to construct a countably infinite family of connected graphs whose capacity is strictly larger than the independence number of any finite (strong) power of the graph.
@@
Therefore, it follows that for all $k\in\mathbb{N}$, $\Theta(\mathsf{G})^{k}\neq\alpha(\mathsf{G}^{k})$ since the independence number is an integer. Consequently, (134) holds.
@@
Next, we consider three approaches to the construction of families of graphs whose Shannon capacities satisfy the condition in (134). The first approach is due to Guo and Watanabe [10] with a simplified proof (the original proof from [10] is provided in Appendix D), and the second and third approaches of constructing such graph families are original.
@@ 5.1 First approach
where the last equality holds by Lemma 2.2. Next, by the assumption that $\Theta(\mathsf{H})>\alpha(\mathsf{H})$ and since $\Theta(\mathsf{H}^{m})=\Theta(\mathsf{H})^{m}\geq\alpha(\mathsf{H})^{m}$ for all $m\in\mathbb{N}$ with strict inequality if $m=1$ (by assumption), it follows that for all $k\in\mathbb{N}$,
@@
$=\bigl(\Theta(\mathsf{G})+\Theta(\mathsf{H})\bigr)^{k}.$ (140)
@@
$\Theta(\operatorname{\mathsf{C}}_{2k+1})\geq\sqrt{\alpha(\operatorname{\mathsf{C}}_{2k+1}^{2})}=\sqrt{k^{2}+\bigg\lfloor\frac{k}{2}\bigg\rfloor}>k=\alpha(\operatorname{\mathsf{C}}_{2k+1}).$ (144)
@@ 5.2 Second approach
Let $m\in\mathbb{N}$ be such that $\sqrt{m}\in\mathbb{Q}$. Let $n_{0}\in\mathbb{N}$ be the smallest natural number such that $n_{0}\sqrt{m}\in\mathbb{N}$. Suppose that $\sqrt{m}\notin\mathbb{N}$. Then, there exists $\ell\in\mathbb{N}$ such that $0<\sqrt{m}-\ell<1$. Define $n_{1}=n_{0}(\sqrt{m}-\ell)$. Since $n_{0}\sqrt{m}\in\mathbb{N}$ and $n_{0}\ell\in\mathbb{N}$, it follows that $n_{1}\in\mathbb{N}$, and also $0<n_{1}<n_{0}$. In addition, $n_{1}\sqrt{m}=n_{0}m-\ell(n_{0}\sqrt{m})\in\mathbb{N}$, which contradicts the minimality of $n_{0}$. Hence, $\sqrt{n}\in\mathbb{N}$. ∎
@@
Let $r\geq 2$, and let $\mathsf{G}_{1},\mathsf{G}_{2},\ldots,\mathsf{G}_{r}$ be graphs such that for every $\ell\in[r]$:
@@
    The graph $\mathsf{G}_{\ell}$ is either a Kneser graph, a self-complementary strongly regular graph, or a self-complementary vertex-transitive graph.
@@
    There exists a single $\ell_{0}\in[r]$ such that $\mathsf{G}_{\ell_{0}}$ is either self-complementary vertex-transitive or self-complementary strongly regular on $n_{\ell_{0}}$ vertices, where $n_{\ell_{0}}$ is not the square of an integer.
Let $p\in\mathbb{N}[x_{1},\ldots,x_{r}]$ be a polynomial with nonzero coefficients in $\mathbb{N}$ such that the following hold:
@@
    The variable $x_{\ell_{0}}$ does not occur in at least one monomial of $p(x_{1},\ldots,x_{r})$.
@@
For all $\ell\in[r]$, $\Theta(\mathsf{G}_{\ell})=\vartheta(\mathsf{G}_{\ell})$ (this follows from the first assumption on $\mathsf{G}_{\ell}$). Thus, by Corollary 3.1,
@@
Let $r\geq 2$, and let $\mathsf{G}_{1},\mathsf{G}_{2},\ldots,\mathsf{G}_{r}$ be graphs such that for every $\ell\in[r]$:
@@
Next, we present an additional example based on Paley graphs.
@@
is not attained by the independence number of any finite strong power of $\mathsf{G}$. This follows from Corollary 5.2, since the Paley graphs are self-complementary and strongly regular, and the restrictions on $q_{1},\ldots,q_{\ell}$ guarantee that the conditions of Corollary 5.2 hold.